Differentiate between enthalpy driven and entropy driven reactions with the help of
suitable examples.
1
Expert's answer
2020-02-17T07:10:40-0500

Solution.

Entropy.

For a spontaneous process (chemical reaction), the change in entropy is always greater than 0. For example:

"Cu + 2AgNO3 = 2Ag + Cu(NO3)2"

If the entropy change is less than 0, then the reaction goes in the opposite direction (or you need to apply additional energy for this process to proceed).

Enthalpy.

If the change in the enthalpy of the reaction is less than 0, then this reaction proceeds with the release of heat. For example:

"S + O2 = SO2"

If the change in the enthalpy of a reaction is greater than 0, then this reaction proceeds with the absorption of heat. For example:

"CaCO3 = CaO + CO2"

Answer:

Entropy.

For a spontaneous process (chemical reaction), the change in entropy is always greater than 0. For example:

"Cu + 2AgNO3 = 2Ag + Cu(NO3)2"

If the entropy change is less than 0, then the reaction goes in the opposite direction (or you need to apply additional energy for this process to proceed).

Enthalpy.

If the change in the enthalpy of the reaction is less than 0, then this reaction proceeds with the release of heat. For example:

"S + O2 = SO2"

If the change in the enthalpy of a reaction is greater than 0, then this reaction proceeds with the absorption of heat. For example:

"CaCO3 = CaO + CO2"